OpenStudy (anonymous):

rocket-powered aleds are used to test human response to acceleration. if a rocket-powered sled accelerates to a speed of 444 m/s in 1.8 seconds, what is its acceleration and how far does it travel?

OpenStudy (anonymous):

1. given \[v = a . t\] and at \[t = 1.8s : v = 444 m/s \] thus is the acceleration \[ a = 246.67 m/s^2\] 2. how far: \[y = 1 / 2 a t ^ 2 = 399.6 m\]